# AllSpark: a multimodal spatiotemporal general model

### 31 Dec 2023 v1

@@ 6 Conclusion
Leveraging multimodal data is an inherent requirement for intelligent models to achieve geographic object cognition. Inspired by human cognitive systems and linguistic philosophy, we propose that the construction of multimodal models should follow the fundamental principle of Language as Reference Framework (LaRF). Guided by this principle, we use language to balance the cohesion and autonomy of modalities, presenting a unified intelligent model, AllSpark, encompassing 13 spatiotemporal modalities. Experimental results demonstrate that AllSpark exhibits excellent adaptability and application potential across various spatiotemporal modalities, highlighting the feasibility and potential of constructing multimodal models with LaRF. AllSpark remains an initial exploratory work, and in the future, we aim to delve deeper into mechanisms guided by natural language, efficient generalization of large models, and the transition to an interactive paradigm.

### 25 Feb 2024 v2

@@ VI Conclusion
Leveraging multimodal data is an inherent requirement for intelligent models to achieve geographic object cognition. Inspired by human cognitive systems and linguistic philosophy, we propose that the construction of multimodal models follow the fundamental principle of Language as Reference Framework (LaRF). Guided by this principle, we use language to balance the cohesion and autonomy of modalities, presenting a unified intelligent model, AllSpark, encompassing thirteen spatio-temporal modalities. The experimental results demonstrated that AllSpark exhibited excellent adaptability and application potential across various spatio-temporal modalities, highlighting the feasibility and potential of constructing multimodal models with LaRF. AllSpark remains an initial exploratory work, and in the future, we aim to delve deeper into the mechanisms guided by natural language, the efficient generalization of large models, and the transition to an interactive paradigm.
TABLE XVI: List of prompts
Dimension
Modal Prompt
1D Text 1.Please determine if this movie review is positive or negative?
Code -
2D RGB 1.This remote sensing image belongs to which of the following categories: [Category list of the NWPU dataset] 2.Describe this remote sensing image briefly 3.Find a word that is most relevant to this remote sensing image 4.Describe the key elements in this remote sensing image
MSI 1.Based on the multi-spectral imagery feature description, please classify this object 2.Given the following multi-spectral imagery characteristics, please output the most fitting scene label
HSI 1.Given the spectral information, can you help determine which class this pixel belongs to? 2.Here is the spectral data for a pixel. Considering the typical characteristics of land cover classes, could you provide a detailed analysis and suggest the most likely class for this pixel? 3.The spectral information for a pixel is given, but the data is noisy. Given the potential variability, which land cover classes should be considered as possible candidates for this pixel?
Table 1.Please utilize the provided air quality indicators to accurately predict the concentrations of PM2.5 in the atmosphere 2.Given the air quality indicators, please provide a prediction for the PM2.5 levels at this particular moment
Oblique Photography 1.Please reconstruct the 3D model from the given images of five different views.
Trajectory 1.Based on their past positions and movements in a crowded environment, predict the future trajectory of a selected pedestrian. 2.Using the pedestrian trajectory data, along with additional information about the surrounding environment, predict the future path of the pedestrian. 3.Given the current and past positions of a pedestrian and their neighboring pedestrians, predict the main pedestrian’s trajectory
SAR 1.Based on the SAR imagery feature description, please classify this object 2.Given the following SAR imagery characteristics, please output the most fitting scene label
Infrared 1.Given an infrared image of a person, find and highlight the same person in a set of visible light images, paying close attention to features like clothing, posture, and gait 2.Analyze the characteristics of a person in a visible light image, such as clothing texture, color, and body shape. Then, locate the person with matching features in a series of infrared images
Graph 1.Given the current traffic data including vehicle flow rate, average speed, and time of day from the METRLA dataset, predict the traffic flow 2.Analyze the historical data on vehicle speeds and flow rates from the METRLA dataset for the past week. Identify any patterns or trends and predict the traffic conditions
3D Video 1.Please analyze this video and describe the characteristics of the action in detail 2.Watch this short clip and provide a step-by-step description of the action sequence
Point Cloud 1.Classify the provided point cloud sample into the correct category 2.Look at the point cloud data characteristics and classify the object 3.Please analyze the given point cloud dataset and determine which category it belongs to. Focus on the shape and structure evident in the point cloud